# Cohort-Based Active Modality Acquisition

### v2

@@ Abstract
Real-world machine learning applications often involve data from multiple modalities that must be integrated effectively to make robust predictions. However, in many practical settings, not all modalities are available for every sample, and acquiring additional modalities can be costly. This raises the question: which samples should be prioritized for additional modality acquisition when resources are limited? While prior work has explored individual-level acquisition strategies and training-time active learning paradigms, test-time and cohort-based acquisition remain underexplored. We introduce Cohort-based Active Modality Acquisition (CAMA), a novel test-time setting to formalize the challenge of selecting which samples should receive additional modalities. We derive acquisition strategies that leverage a combination of generative imputation and discriminative modeling to estimate the expected benefit of acquiring missing modalities based on common evaluation metrics. We also introduce upper-bound heuristics that provide performance ceilings to benchmark acquisition strategies. Experiments on multimodal datasets with up to 15 modalities demonstrate that our proposed imputation-based strategies can more effectively guide the acquisition of additional modalities for selected samples compared with methods relying solely on unimodal information, entropy-based guidance, or random selection. We showcase the real-world relevance and scalability of our method by demonstrating its ability to effectively guide the costly acquisition of proteomics data for disease prediction in a large prospective cohort, the UK Biobank (UKBB). Our work provides an effective approach for optimizing modality acquisition at the cohort level, enabling more effective use of resources in constrained settings. [^1] Code will be published on GitHub.
@@ 1 Introduction
Consider a clinical healthcare setting where all patients in a cohort undergo a standard, inexpensive set of initial examinations, such as basic blood tests and anamnesis. However, a more advanced, expensive, or invasive procedure, like genomic sequencing or specialized imaging, could offer crucial diagnostic or prognostic information for a subset of these patients (Huang et al., 2021). Given a limited budget or capacity for the more advanced procedure, the central question becomes: which patients should receive these additional resources to maximize the overall diagnostic yield or improve treatment outcomes across the entire cohort?
@@
Figure 1: Motivational example for CAMA determining the added value of obtaining the MRI modality. (A) A heterogeneous cohort for which each sample has $P$ distinct modalities. (B) Instead of using the initial subset logit scores ${\textnormal{s}}_{i}^{\textnormal{avail}}$, a generative model $f_{\textnormal{imp}}$ imputes the target missing modality for every patient in the cohort. This yields imputed, augmented-modality logit scores $\{{\textnormal{s}}_{i,k}^{\textnormal{imp}}\}_{k=1}^{K}$ that approximate the logits as if that modality were available. When exactly one modality is missing, these scores approximate ${\textnormal{s}}_{i}^{\textnormal{full}}$; otherwise they estimate the counterfactual with only the imputed modality added. (C) An AF utilizes these scores to rank samples by acquisition priority. The graph demonstrates how the global performance metric improves from the initial baseline towards the performance of a model with access to complete multimodal data, as an increasing fraction of the cohort receives the additional modality. This acquisition process is guided by the proposed strategies operating under the acquisition budget constraint $\beta$.
@@
Balancing potential gains from data modalities against the costs and complexities of acquisition is not unique to healthcare. In remote sensing, for instance, decisions must be made regarding which geographical areas warrant costly high-resolution satellite imagery to supplement widely available, lower-resolution data, aiming to optimize regional environmental monitoring under budget constraints. Likewise, in industrial quality assurance, manufacturers could decide which components from a production batch should undergo detailed, time-consuming testing in addition to rapid, standard visual inspections to effectively identify defects at a batch level. The topic of efficient data acquisition has led to several established paradigms in machine learning, such as Active Learning (AL) (Holzmüller et al., 2023), Active Feature Acquisition (AFA) (Shim et al., 2018), Active Modality Acquisition (AMA) (Kossen et al., 2023), and multimodal learning with missing data (Wu et al., 2024). However, previous research predominantly centers on optimizing acquisition for individual samples and often does not directly address test-time budget constraints for an entire cohort. Consequently, the strategic, test-time acquisition of entire modalities from a cohort perspective remains a significant, largely unaddressed gap. This setting involves deciding, for a given batch of new samples where different subsets of modalities are available, which specific samples should receive an additional, costly modality to best achieve a global objective, e.g., maximizing overall predictive performance or diagnostic accuracy for the cohort, subject to budget constraints. We hypothesize that imputation-based acquisition functions (AFs) can effectively guide resource allocation under cohort-level constraints. The main contributions of this work are as follows:
@@
  • Development of AFs for CAMA We propose a theoretical framework, derived from established evaluation metrics, e.g., Area Under the Receiver Operating Characteristic (AUROC) and Area Under the Precision-Recall Curve (AUPRC), that provides a foundation for developing AFs within the CAMA setting.
@@ 2 Related Work
AL seeks to enhance model training by selecting unlabeled data points for annotation by an oracle (Settles, 2012; Ren et al., 2022; Li et al., 2025). Our methodology draws significantly from AL principles, particularly in the development of an AF to guide the selection process. Consequently, established AL strategies and concepts, such as those rooted in measuring uncertainty (Settles, 2012; Han & Kang, 2021; Hoarau et al., 2025; Raj & Bach, 2022; Ma et al., 2019) or using generative models (Tran et al., 2019; Zhu & Bento, 2017; Zhang et al., 2024; Ma et al., 2019; Peis et al., 2022), are central to our work. Existing work on multimodal acquisition (Rudovic et al., 2019; Das et al., 2022), batch-level selection (Ash et al., 2020; Kirsch et al., 2019; Holzmüller et al., 2023), and balanced AL (Aggarwal et al., 2020; Shen et al., 2023; Zhang et al., 2023; Hoarau et al., 2025) is especially relevant. Our approach, however, diverges from the conventional goals of directly optimizing model training or seeking labels for specific data points: We aim to identify those samples for which the acquisition of an additional data modality would be most beneficial.
@@
AFA builds upon AL by focusing on selecting the most informative individual features for a given sample, often considering their acquisition costs (Rahbar et al., 2025). Similar to AL approaches, methods for AFA encompass a diverse range of techniques, including strategies based on measuring uncertainty (Hoarau et al., 2025; Astorga et al., 2024), the use of generative models (Li & Oliva, 2021; 2024; Gong et al., 2019; Zannone et al., 2019), and Reinforcement Learning (RL) (Valancius et al., 2024; Janisch et al., 2020; Kleist et al., 2025; Shim et al., 2018; Baja et al., 2025). Other common methodologies involve batch-level perspectives (Asgaonkar et al., 2024), leveraging information bottlenecks (Norcliffe et al., 2025), or employing the Kullback-Leibler Divergence (KL-Divergence) (Natarajan et al., 2018). Some AFA techniques rely on gradient calculations (Ghosh & Lan, 2023), while distinct approaches are formulated as individual, sequential recommender systems (Freyberg et al., 2024; Vivar et al., 2020). At an application level, even Large Language Models (LLMs), such as Med-PaLM 2 (Singhal et al., 2025), could be employed for AFA, although such deployments remain unexplored in this context. While our setting shares the core idea of AFA, it differs significantly: We are not concerned with the selection of individual features, but rather with identifying which entire data modalities to acquire. Furthermore, this decision-making process is applied at the cohort level, rather than optimizing for individual samples.
@@
AMA can be conceptualized as an extension of AFA, distinguished by its focus on selecting entire data modalities rather than individual features or labels. Prominent related research includes approaches employing RL for multimodal temporal data (Kossen et al., 2023) and methods utilizing submodular optimization in conjunction with Shapley values (Shapley, 1953; He et al., 2024). The approach by Kossen et al. (2023) differs from ours through its reliance on RL, whereas He et al. (2024) primarily investigate how modalities affect optimal learning performance. Further studies have explored the use of Gaussian mixtures within Bayesian optimal experimental design to enhance data acquisition efficiency for model training (Long, 2022). This objective differs from ours, as our focus is not on improving the model training process itself, but rather on optimizing performance for a downstream task at test time. The relative sparsity of existing work for AMA underscores the significance of the research gap that our proposed setting, i.e., CAMA  aims to address.
@@
Research in multimodal learning with missing data modalities offers techniques for robustly handling incomplete datasets. These methods are broadly classified into strategy design aspects, i.e., architecture-focused designs and model combinations, and data processing aspects, i.e., representation learning and modality imputation (Wu et al., 2024). Acknowledging the utility of these approaches, our work emphasizes imputation-based strategies, and thus this paragraph highlights those methods. Imputation of missing features is commonly performed using Auto Encoders (AEs) (Hinton & Zemel, 1993), Variational Auto Encoders (VAEs) (Kingma & Welling, 2014), Generative Adversarial Networks (GANs) (Goodfellow et al., 2014), or Denoising Diffusion Probabilistic Models (DDPMs) (Ho et al., 2020; Rombach et al., 2022). These methods naturally extend to multiple modalities, for example, with VAE-based (Wesego & Rooshenas, 2024; Sutter et al., 2021; Lewis et al., 2021) and DDPM-based (Wang et al., 2023) approaches. Notably, the latter, i.e., IMDer (Wang et al., 2023), a multimodal deep learning architecture that imputes missing values with DDPMs in latent spaces, is adapted in our work (Section˜5). However, this research area focuses on handling absent modalities rather than deciding which ones to acquire.
@@ 3 Problem Formulation
  • ${\textnormal{s}}_{i}^{\textnormal{full}}$: The full score, a theoretical oracle score that would be computed if all modalities were available. This is unknown for samples with missing data modalities.
  • $\{{\textnormal{s}}_{i,k}^{\textnormal{imp}}\}_{k=1}^{K}$: A set of $K$ imputed scores that estimate the unknown ${\textnormal{s}}_{i}^{\textnormal{full}}$ using only the available data modalities.
For instance, given the example from Figure˜1, in a simple clinical setting with a cheap, universally available base modality, e.g., cardiac biomarkers such as troponin or B-type natriuretic peptide (BNP), and an expensive additional modality, e.g., cardiac MRI, ${\textnormal{s}}_{i}^{\textnormal{avail}}$ would be the score from the blood tests alone, while ${\textnormal{s}}_{i}^{\textnormal{full}}$ would be the score using both tests and MRI. To compute these scores, we assume a single model $f$ parameterized by $\bm{\theta}$} that can process any subset of modalities. The available and full scores are thus:
@@
${\textnormal{s}}_{i}^{\textnormal{full}}$ $=f({\bm{x}}_{i}^{\textnormal{full}},{\bm{\theta}})$ (2)
where ${\bm{x}}_{i}^{\textnormal{avail}}$ and ${\bm{x}}_{i}^{\textnormal{full}}$ represent the feature sets for the available and full modalities, respectively. To estimate the full score without costly acquisition, we use a generative imputation model $f_{\textnormal{imp}}$. This model generates a set of $K$ plausible embeddings that enable the classifier $f$ to predict the scores $\{{\textnormal{s}}_{i,k}^{\textnormal{imp}}\}_{k=1}^{K}$. These imputation-based scores form the basis of our acquisition functions.
@@
${\textnormal{s}}_{i}({\mathcal{S}})=\begin{cases}{\textnormal{s}}_{i}^{\textnormal{full}}&\text{if }i\in{\mathcal{S}}\\ {\textnormal{s}}_{i}^{\textnormal{avail}}&\text{if }i\notin{\mathcal{S}}\end{cases}.$ (3)
@@ 4 Acquisition Function Strategies
  • Baseline Information Strategies: These strategies make decisions using only information from the initially available modality, i.e., without any imputation, such as its predicted uncertainty or probability.
@@
Table 2: Summary of AF Strategies.
Category Strategies Input Variables Ranking Criteria
Oracle AUROC, AUPRC True labels & full scores Greedy selection for maximum gain.
Upper-Bound KL-Divergence, Rank, Uncertainty True full scores True change in prediction, cohort rank, or uncertainty.
Imputation-Based KL-Divergence, Rank, Uncertainty & Probability Imputed full scores Expected change in prediction, rank, or uncertainty.
Baselines Uncertainty, Probability Unimodal scores Uncertainty or probability using the available modality.
Random Random None Random selection.
For evaluation, we introduce a metric that describes the cumulative performance of an AF, normalized by the total possible gain achievable by transitioning all samples to full multimodal performance (see Figure˜1 C, for an illustrative curve). Let $M_{\text{AF}}(b)$ denote the performance curve of an AF strategy for a primary metric $M$ as a function of the budget fraction $b$ of the acquisition budget $\beta$, $M_{\text{uni}}$ the performance of the unimodal baseline, and $M_{\text{multi}}$ the performance of the full multimodal model. The normalized area of gain for an acquisition function AF, which measures the portion of achievable performance gain captured across different budgets, is defined as the area under the performance gain curve, normalized by the maximum possible gain (Equation˜5). Intuitively, a value of 0 indicates no improvement over the unimodal baseline across budgets. A value of $1$ indicates matching the full multimodal performance on average across budgets. Values greater than $1$ occur when the cohort’s performance at intermediate budgets temporarily exceeds the multimodal cohort.
$G_{\text{full}}^{M}(\text{AF})=\frac{\int_{0}^{1}(M_{\text{AF}}(b)-M_{\text{uni}})\,db}{M_{\text{multi}}-M_{\text{uni}}}$ (5)

### v3

@@ Abstract
Real-world machine learning applications often involve data from multiple modalities that must be integrated effectively to make robust predictions. However, in many practical settings, not all modalities are available for every sample, and acquiring additional modalities can be costly. This raises the question: which samples should be prioritized for additional modality acquisition when resources are limited? While prior work has explored individual-level acquisition strategies and training-time active learning paradigms, test-time and cohort-based acquisition remain underexplored. We introduce Cohort-based Active Modality Acquisition (CAMA), a novel test-time setting to formalize the challenge of selecting which samples should receive an additional modality. We derive acquisition strategies that leverage a combination of generative imputation and discriminative modeling to estimate the expected benefit of acquiring a missing modality based on common evaluation metrics. We also introduce upper-bound heuristics that provide performance ceilings to benchmark acquisition strategies. Experiments on multimodal datasets with up to 15 modalities demonstrate that our proposed imputation-based strategies can more effectively guide the acquisition of an additional modality for selected samples compared with methods relying solely on pre-acquisition information, entropy-based guidance, or random selection. We showcase the real-world relevance and scalability of our method by demonstrating its ability to effectively guide the costly acquisition of proteomics data for disease prediction in a large prospective cohort, the UK Biobank (UKBB). Our work provides an effective approach for optimizing modality acquisition at the cohort level, enabling more effective use of resources in constrained settings. [^1] Code will be published on GitHub.
@@ 1 Introduction
Consider a clinical healthcare setting where all patients in a cohort undergo a standard, inexpensive set of initial examinations, such as basic blood tests and anamnesis. However, a more advanced, expensive, or invasive procedure, like genomic sequencing or specialized imaging, could offer crucial diagnostic or prognostic information for a subset of these patients (huang_what_2021). Given a limited budget or capacity for the more advanced procedure, the central question becomes: which patients should receive this additional resource to maximize the overall diagnostic yield or improve treatment outcomes across the entire cohort?
@@
Figure 1: Motivational example for CAMA determining the added value of obtaining the MRI modality. (A) A heterogeneous cohort for which each sample has $P$ distinct modalities. (B) Instead of using the initial subset logit scores ${\textnormal{s}}_{i}^{\textnormal{avail}}$, a generative model $f_{\textnormal{imp}}$ imputes the target missing modality for every patient in the cohort. This yields imputed, augmented-modality logit scores $\{{\textnormal{s}}_{i,k}^{\textnormal{imp}}\}_{k=1}^{K}$ that approximate the logits as if that modality were available. These scores approximate ${\textnormal{s}}_{i}^{\textnormal{acquired}}$, i.e., the counterfactual with only the imputed modality added. (C) An AF utilizes these scores to rank samples by acquisition priority. The graph demonstrates how the global performance metric improves from the initial baseline towards the performance of a model with access to post-acquisition data, as an increasing fraction of the cohort receives the additional modality. This acquisition process is guided by the proposed strategies operating under the acquisition budget constraint $\beta$.
@@
For healthcare, budgets are often resource-specific rather than flexible. For example, a hospital may have a fixed capacity for one MRI scanner, or a cohort may have a specific grant for one modality. The critical decision is prioritizing access to that single resource across the cohort, and not necessarily dynamically acquiring for different modalities per patient. Consider a healthcare system that can afford 1,000 expensive tests for a 100,000-person cohort. The goal is to improve health outcomes across the whole population, and this typically happens through resource allocation: who receives preventive interventions, who gets enrolled in clinical trials, who is flagged for closer monitoring. These decisions depend on accurate risk stratification. A global ranking of all 100,000 individuals by predicted risk (measured, for example, by Area Under the Receiver Operating Characteristic (AUROC)) becomes the tool through which such allocation decisions are made. The question is: which 1,000 patients should we test so that our final ranking of all 100,000 is as accurate as possible? Balancing potential gains from data modalities against the costs and complexities of acquisition is not unique to healthcare. In remote sensing, for instance, decisions must be made regarding which geographical areas warrant costly high-resolution satellite imagery to supplement widely available, lower-resolution data, aiming to optimize regional environmental monitoring under budget constraints. Likewise, in industrial quality assurance, manufacturers could decide which components from a production batch should undergo detailed, time-consuming testing in addition to rapid, standard visual inspections to effectively identify defects at a batch level. The topic of efficient data acquisition has led to several established paradigms in machine learning, such as Active Learning (AL) (holzmuller_framework_2023), Active Feature Acquisition (AFA) (shim_joint_2018), Active Modality Acquisition (AMA) (kossen_active_2023), and multimodal learning with missing data (wu_deep_2024). However, previous research predominantly centers on optimizing acquisition for individual samples and often does not directly address test-time budget constraints for an entire cohort. Consequently, the strategic, test-time acquisition of an additional modality from a cohort perspective remains a significant, largely unaddressed gap. This setting involves deciding, for a given batch of new samples where different subsets of modalities are available, which specific samples should receive an additional, costly modality to best achieve a global objective, e.g., maximizing overall predictive performance or diagnostic accuracy for the cohort, subject to budget constraints. We hypothesize that imputation-based acquisition functions (AFs) can effectively guide resource allocation under cohort-level constraints. The main contributions of this work are as follows:
@@
  • Development of AFs for CAMA We propose a theoretical framework, derived from established evaluation metrics, e.g., AUROC and Area Under the Precision-Recall Curve (AUPRC), that provides a foundation for developing AFs within the CAMA setting.
@@ 2 Related Work
AL seeks to enhance model training by selecting unlabeled data points for annotation by an oracle (settles_active_2012; ren_survey_2022; li_survey_2025). Our methodology draws significantly from AL principles, particularly in the development of an AF to guide the selection process. Consequently, established AL strategies and concepts, such as those rooted in measuring uncertainty (settles_active_2012; han_active_2021; hoarau_reducing_2025; raj_convergence_2022; ma_eddi_2019) or using generative models (tran_bayesian_2019; zhu_generative_2017; zhang_generative_2024; ma_eddi_2019; peis_missing_2022), are central to our work. Existing work on multimodal acquisition (rudovic_multi-modal_2019; das_mavic_2022), batch-level selection (ash_deep_2020; kirsch_batchbald_2019; holzmuller_framework_2023), and balanced AL (aggarwal_active_2020; shen_towards_2023; zhang_algorithm_2023; hoarau_reducing_2025) is especially relevant. Our approach, however, diverges from the conventional goals of directly optimizing model training or seeking labels for specific data points: We aim to identify those samples for which the acquisition of an additional data modality would be most beneficial.
@@
AFA builds upon AL by focusing on selecting the most informative individual features for a given sample, often considering their acquisition costs (rahbar_survey_2025). Similar to AL approaches, methods for AFA encompass a diverse range of techniques, including strategies based on measuring uncertainty (hoarau_reducing_2025; astorga_active_2024), the use of generative models (li_active_2021; li_distribution_2024; gong_icebreaker_2019; zannone_odin_2019), and Reinforcement Learning (RL) (valancius_acquisition_2024; janisch_classification_2020; kleist_evaluation_2025; shim_joint_2018; baja_measure_2025). Other common methodologies involve batch-level perspectives (asgaonkar_generator_2024), leveraging information bottlenecks (norcliffe_information_2025), or employing the Kullback-Leibler Divergence (KL-Divergence) (natarajan_whom_2018). Some AFA techniques rely on gradient calculations (ghosh_difa_2023), while distinct approaches are formulated as individual, sequential recommender systems (freyberg_mint_2024; martel_peri-diagnostic_2020). At an application level, even Large Language Models (LLMs), such as Med-PaLM 2 (singhal_toward_2025), could be employed for AFA, although such deployments remain unexplored in this context. While our setting shares the core idea of AFA, it differs significantly: We are not concerned with the selection of individual features, but rather with identifying which entire data modalities to acquire. Furthermore, this decision-making process is applied at the cohort level, rather than optimizing for individual samples.
@@
AMA can be conceptualized as an extension of AFA, distinguished by its focus on selecting entire data modalities rather than individual features or labels. Prominent related research includes approaches employing RL for multimodal data (kossen_active_2023; jain_test-time_2025; li_towards_2025) and methods utilizing submodular optimization in conjunction with Shapley values (kuhn_value_1953; he_efficient_2024). The approach by kossen_active_2023 differs from ours through its reliance on RL, whereas he_efficient_2024 primarily investigate how modalities affect optimal learning performance. Further studies have explored the use of Gaussian mixtures within Bayesian optimal experimental design to enhance data acquisition efficiency for model training (long_multimodal_2022). This objective differs from ours, as our focus is not on improving the model training process itself, but rather on optimizing performance for a downstream task at test time. The relative sparsity of existing work for AMA underscores the significance of the research gap that our proposed setting, i.e., CAMA  aims to address.
@@
Research in multimodal learning with missing data modalities offers techniques for robustly handling incomplete datasets. These methods are broadly classified into strategy design aspects, i.e., architecture-focused designs and model combinations, and data processing aspects, i.e., representation learning and modality imputation (wu_deep_2024). Acknowledging the utility of these approaches, our work emphasizes imputation-based strategies, and thus this paragraph highlights those methods. Imputation of missing features is commonly performed using Auto Encoders (AEs) (hinton_autoencoders_1993), Variational Auto Encoders (VAEs) (kingma_auto-encoding_2014), Generative Adversarial Networks (GANs) (goodfellow_generative_2014), or Denoising Diffusion Probabilistic Models (DDPMs) (ho_denoising_2020; rombach_high-resolution_2022). These methods naturally extend to multiple modalities, for example, with VAE-based (wesego_score-based_2024; sutter_generalized_2021; lewis_accurate_2021) and DDPM-based (wang_incomplete_2023) approaches. Notably, the latter, i.e., IMDer (wang_incomplete_2023), a multimodal deep learning architecture that imputes missing values with DDPMs in latent spaces, is adapted in our work (Section˜5). However, this research area focuses on handling absent modalities rather than deciding which ones to acquire.
@@ 3 Problem Formulation
  • ${\textnormal{s}}^{\text{acquired}}_{i}$: The acquired score, computed using the sample’s available modalities plus the newly acquired modality.
  • $\{{\textnormal{s}}_{i,k}^{\textnormal{imp}}\}_{k=1}^{K}$: A set of $K$ imputed scores that estimate the unknown ${\textnormal{s}}_{i}^{\textnormal{acquired}}$ using only the available data modalities.
For instance, given the example from Figure˜1, in a simple clinical setting with a cheap, universally available base modality, e.g., cardiac biomarkers such as troponin or B-type natriuretic peptide (BNP), and an expensive additional modality, e.g., cardiac MRI, ${\textnormal{s}}_{i}^{\textnormal{avail}}$ would be the score from the blood tests alone, while ${\textnormal{s}}_{i}^{\textnormal{acquired}}$ would be the score using both tests and MRI. To compute these scores, we assume a single model $f$ parameterized by $\bm{\theta}$} that can process any subset of modalities. The available and acquired scores are thus:
@@
${\textnormal{s}}_{i}^{\textnormal{acquired}}$ $=f({\bm{x}}_{i}^{\textnormal{acquired}},{\bm{\theta}})$ (2)
where ${\bm{x}}_{i}^{\textnormal{avail}}$ and ${\bm{x}}_{i}^{\textnormal{acquired}}$ represent the feature sets for the available and acquired modalities, respectively. To estimate the acquired score without costly acquisition, we use a generative imputation model $f_{\textnormal{imp}}$. This model generates a set of $K$ plausible embeddings that enable the classifier $f_{C}$ to predict the scores $\{{\textnormal{s}}_{i,k}^{\textnormal{imp}}\}_{k=1}^{K}$. These imputation-based scores form the basis of our acquisition functions.
@@
${\textnormal{s}}_{i}({\mathcal{S}})=\begin{cases}{\textnormal{s}}_{i}^{\textnormal{acquired}}&\text{if }i\in{\mathcal{S}}\\ {\textnormal{s}}_{i}^{\textnormal{avail}}&\text{if }i\notin{\mathcal{S}}\end{cases}.$ (3)
@@
Table 2: Summary of AF Strategies.
Category Strategies Input Variables Ranking Criteria
Oracle AUROC, AUPRC True labels & acquired scores Greedy selection for maximum gain.
Upper-Bound KL-Divergence, Rank, Uncertainty True acquired scores True change in prediction, cohort rank, or uncertainty.
Imputation-Based KL-Divergence, Rank, Uncertainty & Probability Imputed acquired scores Expected change in prediction, rank, or uncertainty.
Baselines Uncertainty, Probability Pre-acquisition scores Uncertainty or probability using the available modality.
Random Random None Random selection.
@@ 4 Acquisition Function Strategies
  • Baseline Information Strategies: These strategies make decisions using only information from the initially available modalities, i.e., without any imputation and pre-acquisition, such as its predicted uncertainty or probability.
@@
Intuitively, these acquisition functions approximate the expected information gain (EIG) from acquiring an additional modality. In our setting, EIG quantifies the expected improvement in a chosen performance metric given the additional information that would become available through a new modality.
For evaluation, we introduce a metric that describes the cumulative performance of an AF, normalized by the total possible gain achievable by transitioning all samples to post-acquisition performance (see Figure˜1 C, for an illustrative curve). Let $M_{\text{AF}}(b)$ denote the performance curve of an AF strategy for a primary metric $M$ as a function of the budget fraction $b$ of the acquisition budget $\beta$, $M_{\text{pre}}$ the performance of the pre-acquisition baseline, and $M_{\text{post}}$ the performance of the post-acquisition model. The normalized area of gain for an acquisition function AF, which measures the portion of achievable performance gain captured across different budgets, is defined as the area under the performance gain curve, normalized by the maximum possible gain (Equation˜5). Intuitively, a value of 0 indicates no improvement over the pre-acquisition baseline across budgets. A value of $1$ indicates matching the post-acquisition performance on average across budgets. Values greater than $1$ occur when the cohort’s performance at intermediate budgets temporarily exceeds the post-acquisition cohort as detailed later and shown in Figure˜3.
$G_{\text{full}}^{M}(\text{AF})=\frac{\int_{0}^{1}(M_{\text{AF}}(b)-M_{\text{pre}})\,db}{M_{\text{post}}-M_{\text{pre}}}$ (5)